Grade 9Physical Chemistry

How can grams of kcl are present in 4.3 ×10^24 solve this problem??

Asma
7 Years agoGrade 9

1 Answer

Khimraj
7 Years ago
mole of KCl = 4.3*1024/6.022140857 × 1023
 = 7.14 mole
in gram = 7.14*74.55 = 532.31 gm
Hope it clears......................